Find the volume of the solid bounded by the planes x = 0, y = 0, z = 0, and x + y + z = 6.
1
Expert's answer
2020-10-27T18:43:07-0400

"V=\iint_D zdydx=\iint_D (6-x-y)dydx"

"D:0\leq x\leq 6, 0\leq y\leq 6-x"


"V=\int_0^6 \int_0^{6-x}(6-x-y)dydx =\int _0^6 (6y-xy-\frac{y^2}{2})|_0^{6-x}dx="

"=\int_0 ^6 (6\cdot(6-x)-x\cdot(6-x)-\frac{(6-x)^2}{2})dx=\int_0^6((6-x)^2-\frac{(6-x)^2}{2})dx="

"=\frac{1}{2} \int_0^6 (6-x)^2dx=\frac{1}{2} \int_0^6 (x-6)^2d(x-6)=(\frac{1}{2}\cdot \frac{(x-6)^3}{3} )|_0^6 ="

"=\frac{(6-6)^3}{6}-\frac{(0-6)^3}{6}=36"